# Characterizing rational homogeneous spaces via $\Csuperscript\C\C^{*}$-actions

### v1

@@ Abstract.
Authors supported by the Department of Mathematics of the University of Trento, and partially supported by INdAM–GNSAGA. The authors would like to thank Jarek Wiśniewski for inspirational conversations and useful remarks.
@@ 1. Introduction
One long standing open problem in the framework of Fano manifolds is the Campana–Peternell (CP, for short) Conjecture, that states that the only Fano manifolds with nef tangent bundle are the rational homogeneous varieties. In other words, for a Fano manifold $X$ the nefness assumption on $T_{X}$ should imply the existence of an affine algebraic group acting transitively on $X$. One usually considers the case in which the Picard number $\rho_{X}$ is equal to one, from which the general case should follow. For more information on the Conjecture and further references we refer the reader to the survey [21].
In view of Mori’s proof of the Hartshorne–Fränkel Conjecture –of which the CP Conjecture is a natural generalization– one of the strategies that has been used to prove partial results is the study of rational curves. If the tangent bundle $T_{X}$ is nef, then the deformations of rational curves in $X$ are unobstructed, and one may try to reconstruct the geometry of $X$ out of certain families of rational curves. Of particular interest in this sense are the so called unsplit and dominating families of rational curves, that allow to consider some sort of infinitesimal-to-global arguments; more concretely, many results in the literature show that the homogeneity of a Fano manifold $X$ of Picard number one can be inferred from the homogeneity of the sets of tangent directions to the curves of an unsplit and dominating family of rational curves at points (see, for instance [20, 12, 27]).
@@
In order to prove Theorem 1.1 we first study, in Section 4, the induced ${\mathbb{C}}^{*}$-actions on the beautiful family ${\mathcal{M}}$ of rational curves in $X$ and on the universal family ${\mathcal{U}}$, without any assumptions on the extremal fixed point components. We prove that both the induced actions on ${\mathcal{M}}$ and ${\mathcal{U}}$ are equalized and that the induced action on the variety ${\mathcal{M}}_{y}$ of curves of ${\mathcal{M}}$ passing by a fixed point $y\in X^{{\mathbb{C}}^{*}}$ has criticality at most two.
@@
In Section 6 we first refine the above results in the case in which the action has an isolated extremal fixed point. In particular, if $y$ is a fixed point belonging to the fixed point component joined to the isolated point by curves in ${\mathcal{M}}$, the action on ${\mathcal{M}}_{y}$ has an extremal fixed point, too. Since the criticality of the action on ${\mathcal{M}}_{y}$ is low, these varieties can be classified: Section 3 is devoted to this task (see Theorems 3.6 and 3.9).
The proof of Theorem 1.1 then proceeds by discarding the non-homogeneous possibilities for ${\mathcal{M}}_{y}$ and by showing that, for every $x\in X$ the variety ${\mathcal{M}}_{x}$ is isomorphic to ${\mathcal{M}}_{y}$. The homogeneity of $X$ then follows from [27, Theorem 1.1], and the fact that $X$ is an irreducible Hermitian symmetric space from the classification of equalized ${\mathbb{C}}^{*}$-actions on homogeneous varieties (see [24, Table 2]).
@@ 2.1. ${\mathbb{C}}^{*}$-actions
$TX_{\mid Y}=T^{+}\oplus T^{0}\oplus T^{-},$ (1)
where $T^{+}$, $T^{0}$, $T^{-}$ are the subbundles of $TX_{\mid Y}$ where ${\mathbb{C}}^{*}$ acts with positive, zero or negative weights, respectively. Then, by local linearization, $T^{0}=TY$ and
@@
A straightforward consequence of the above decomposition is that the tangent bundle of a fixed point component is a quotient of the restriction of the tangent bundle of $X$.
@@ 2.2. Białynicki-Birula decomposition
We refer to [6] for a complete account on the Białynicki-Birula decomposition and its applications, and to [4] for the original reference.
Given any set $S\subset X$, we define:
$X^{\pm}(S):=\{x\in X|\,\,x_{\pm}\in S\},\quad B^{\pm}(S):=\overline{X^{\pm}(S)},$ (4)
@@
We will denote by $B^{\pm}(Y)$ the closure of the BB-cells $X^{\pm}(Y)$. Every cell $X^{\pm}(Y)$ comes with a natural morphism onto $Y$, sending every $x\in X^{\pm}(Y)$ to its limiting point $x_{\pm}$. It was shown by Białynicki-Birula that $X^{\pm}(Y)$ is an affine bundle, locally isomorphic to $N^{\pm}(Y)$.
@@ 2.3. Actions on polarized pairs
A ${\mathbb{C}}^{*}$-action on the polarized pair $(X,L)$ is a ${\mathbb{C}}^{*}$-action on $X$, together with the weight map $\mu_{L}$ determined by the normalized linearization of an ample line bundle $L$. In this case the minimum and maximum of weight map $\mu_{L}$ are achieved precisely at the sink and the source of the action. We then denote by
@@
so that $Y_{0},Y_{r}$ are the sink $Y_{-}$ and then source $Y_{+}$ of the action on $X$, respectively.
@@ 2.4. Equalized actions
A ${\mathbb{C}}^{*}$-action is equalized at $Y\in{\mathcal{Y}}$ if for every $x\in\big{(}X^{-}(Y)\cup X^{+}(Y)\big{)}\setminus Y$ the isotropy group of the action at $x$ is trivial. Equivalently (see [25, Lemma 2.1]), ${\mathbb{C}}^{*}$ acts on $N^{+}(Y)$ with all the weights equal to $+1$ and on $N^{-}(Y)$ with all the weights equal to $-1$. The action is called equalized if it is equalized at every $Y\in{\mathcal{Y}}$.
For an equalized action the closure of any $1$-dimensional orbit is a smooth rational curve, whose $L$-degree may be computed in terms of the weights at its extremal points. (Cf. [25, Lemma 2.2], [28, Corollary 3.2]):
@@
Using the above Lemma together with formula (5), we may compute the anticanonical degrees of ${\mathbb{C}}^{*}$-invariant non-fixed curves:
@@
If the action on $X$ is equalized, and $C$ is the closure of an orbit with sink in a fixed point component $Y$ and source in a fixed point component $Y^{\prime}$, then:
@@
Not even the equalization hypothesis is enough to guarantee that the BB-decompositions associated to a ${\mathbb{C}}^{*}$-action on a smooth projective variety are stratifications, see [23, Example 6.1].
@@ 2.5. Geometric quotients
As shown in [23, Section 2.2], $\mathcal{G}\!X(\tau)$ is the same for any $\tau\in(a_{i},a_{i+1})\cap{\mathbb{Q}}$, and so we set:
@@
If the ${\mathbb{C}}^{*}$-action on $X$ is equalized, the geometric quotients $\mathcal{G}\!X_{0,1}$ and $\mathcal{G}\!X_{r-1,r}$ can be described as follows (see [24, Section 2.1.3] for details): the ${\mathbb{C}}^{*}$-action on $X$ extends to an action on the blowup $X^{\flat}$ of $X$ along its sink and source. Then $\mathcal{G}\!X_{0,1}$ and $\mathcal{G}\!X_{r-1,r}$ are isomorphic to the sink and the source of $X^{\flat}$, that is to ${\mathbb{P}}(N_{Y_{0},X}^{\vee})$ and ${\mathbb{P}}(N_{Y_{r},X}^{\vee})$, respectively (see [24, Section 2.1.3]). In particular, in the case of an action of B-type, i.e., when the extremal fixed point components are divisorial (equivalently, if $\nu^{-}(Y_{-})=\nu^{+}(Y_{+})=1$), we have that $\mathcal{G}\!X_{0,1}\simeq\mathcal{G}\!X_{0,0}\simeq Y_{0}$ and $\mathcal{G}\!X_{r-1,r}\simeq\mathcal{G}\!X_{r,r}\simeq Y_{r}$.
The varieties $\mathcal{G}\!X_{i,i+1}$ are birationally equivalent and the natural birational maps among them fit in the following commutative diagram:
@@
The diagonal maps between geometric quotients are contractions, the maps $s$ and $d$ are smooth blow-ups, the horizontal maps are Atiyah flips, and the squares are fiber products (see [30, Theorem 1.9]).
In particular, in the case in which the action is of B-type the diagram above provides a birational map $\psi:Y_{0}\dashrightarrow Y_{r}$.
The varieties $Z_{i,i+2}$ appearing in Diagram (7) have been interpreted in [23] in terms of Chow quotients of birational modifications of $X$ and the centers of the blow-ups $s$ and $d$ have been described. In this paper we will only use them as resolutions of the birational maps $\mathcal{G}\!X_{i,i+1}\dashrightarrow\mathcal{G}\!X_{i+1,i+2}$. More precisely, let us summarize in the next proposition what we need.
@@
The center of the blow-up $s:Z_{i,i+2}\to\mathcal{G}\!X_{i,i+1}$ (resp. $d:Z_{i,i+2}\to\mathcal{G}\!X_{i+1,i+2}$) is the inverse image in $\mathcal{G}\!X_{i,i+1}$ (resp. $\mathcal{G}\!X_{i+1,i+2}$) of the embedding of $Y_{i+1}$ into $\mathcal{G}\!X_{i+1,i+1}$. Moreover, if $\nu^{-}(Y_{i+1})=1$ (resp. if $\nu^{+}(Y_{i+1})=1$) then $s$ (resp. $d$) is an isomorphism.
@@
The first part of the statement follows from [23, Remark 5.10], the second from [23, Proof of Corollary 7.1]. ∎
@@
Assume that we have a B-type action of criticality three with divisorial sink and source $Y_{0},Y_{3}$ and that $\nu^{+}(Y_{1})=\nu^{-}(Y_{2})=1$. Then the birational map $\psi:Y_{0}\dashrightarrow Y_{3}$ is resolved by a smooth blow-up with center isomorphic to $Y_{1}$ and a smooth blow-down with center isomorphic to $Y_{2}$.
@@ 2.6. Action on rational homogeneous varieties
Standard examples of ${\mathbb{C}}^{*}$-actions can be constructed easily on rational homogeneous (RH, for short) varieties, i.e., projective varieties that are quotients of a semisimple group $G$ by a parabolic subgroup $P$. Without loss of generality we assume that $G$ is the adjoint group of a semisimple Lie algebra ${\mathfrak{g}}$. Following [21, Section 2.2], for instance, an RH variety is completely determined by the Dynkin diagram ${\mathcal{D}}$ of $G$ marked on $\rho(G/P)$ nodes; we will then use the notation ${\mathcal{D}}(I)$ to denote the RH variety determined by the Dynkin diagram ${\mathcal{D}}$ marked in the set $I$ of nodes. Table 1 illustrates how this notation works in the case of irreducible Hermitian symmetric spaces; these are the RH varieties $G/P$ of Picard number one such that $P$ acts transitively on the tangent space $T_{G/P,eP}$, which is $P$-equivariantly isomorphic to the quotient of the associated Lie algebras, ${\mathfrak{g}}/{\mathfrak{p}}$.
Table 1. Irreducible Hermitian symmetric spaces.
@@
The ${\mathbb{C}}^{*}$-action determined by $\sigma_{i}$ on the variety ${\mathcal{D}}(I)=G/P$ satisfies that $eP$ lies in the sink of the action. Furthermore it is the isolated sink of the action if and only if $I=\{i\}$.
@@
Table 3. Irreducible Hermitian symmetric spaces and ${\mathbb{C}}^{*}$-actions with isolated sink on them
@@ 2.7. Families of rational curves
Given a smooth complex projective variety $X$, a family of rational curves on $X$ is an irreducible component ${\mathcal{M}}$ of the normalization $\mathop{\rm RatCurves}\nolimits^{n}(X)$ of the scheme $\mathop{\rm RatCurves}\nolimits(X)$. Such a family comes equipped with a smooth ${\mathbb{P}}^{1}$-fibration $p:{\mathcal{U}}\to{\mathcal{M}}$ and an evaluation morphism $q:{\mathcal{U}}\to X$. The family ${\mathcal{M}}$ is called dominating if $q$ is dominant, and unsplit if ${\mathcal{M}}$ is proper. We will denote by ${\mathcal{M}}_{x}$ the fiber of $q$ over the point $x\in X$.
In this paper we will consider the case in which ${\mathcal{M}}$ is a dominating, unsplit family such that $q:{\mathcal{U}}\to X$ is a smooth morphism, and we will simply say that the family is beautiful. Note that in this case the varieties ${\mathcal{U}}$ and ${\mathcal{M}}$ are smooth as well, and
$\dim{\mathcal{M}}_{x}=-K_{X}\cdot C-2,$ (8)
where $C$ is any element of ${\mathcal{M}}$.
@@
By [17, II.Proposition 3.5], the beauty of the family ${\mathcal{M}}$ is equivalent to the following property: $\nu^{*}T_{X}$ is globally generated – equivalently nef – for the normalization $\nu:{\mathbb{P}}^{1}\to X$ of any curve of the family. In particular, $\mathop{\rm RatCurves}\nolimits^{n}(X)$ is smooth at every point of ${\mathcal{M}}$, and every deformation of a curve of ${\mathcal{M}}$ belongs to ${\mathcal{M}}$. Then, given any irreducible subvariety $Y\subset X$, the set of elements of ${\mathcal{M}}$ parametrizing curves of the family contained in $Y$ is a union of irreducible components of $\mathop{\rm RatCurves}\nolimits^{n}(Y)$.
Note that, if moreover $X$ is a Fano manifold (in particular simply connected), then every ${\mathcal{M}}_{x}$ is connected, therefore irreducible (cf. [15, Proposition 1.6]). This is always the case if the Picard number of $X$ is one.
@@
Curves of ${\mathcal{M}}$ meeting $Z$ are parametrized by ${\mathcal{M}}_{Z}:=p(q^{-1}(Z))$, which is irreducible. If $q^{-1}(Z)\cap p^{-1}({\mathcal{M}}_{Z})=q^{-1}(Z)$ every curve of ${\mathcal{M}}$ meeting $Z$ would be contained in $Z$, but this is not possible since $X$ is ${\mathcal{M}}$)-connected. Therefore $q^{-1}(Z)\cap p^{-1}({\mathcal{M}}_{Z})$ is a proper closed subset of $p^{-1}({\mathcal{M}}_{Z})$, and thus it does not contain a general fiber of $p^{-1}({\mathcal{M}}_{Z})\to{\mathcal{M}}_{Z}$. ∎
@@
Let $X$ be a Fano manifold of Picard number one, and $p:{\mathcal{U}}\to{\mathcal{M}}$ be a beautiful family of rational curves in $X$. If the fiber ${\mathcal{M}}_{x}=q^{-1}(x)$ is a rational homogeneous variety for every $x\in X$, then $X$ is rational homogeneous.
@@
${\mathcal{D}}(I)$ ${\mathcal{M}}_{x}$
@@
In the spirit of the above statement, irreducible Hermitian symmetric spaces can be characterized by the homogeneous varieties ${\mathcal{M}}_{x}$ of lines passing by a given point of the variety. Table 4 describes ${\mathcal{M}}_{x}$ for every type of irreducible Hermitian symmetric space. By comparing it with Table 3, we immediately see that in each case ${\mathcal{M}}_{x}$ coincides with the component $Y_{1}$ of fixed points of weight one.

### v2

@@ Abstract.
Authors supported by the Department of Mathematics of the University of Trento, and partially supported by INdAM–GNSAGA. The authors would like to thank Jarek Wiśniewski for inspirational conversations and useful remarks, and an anonymous referee for his comments and suggestions, which helped improving the exposition of the paper.
@@ 1. Introduction
One long standing open problem in the framework of Fano manifolds is the Campana–Peternell (CP, for short) Conjecture, that states that the only Fano manifolds with nef tangent bundles are rational homogeneous varieties. In other words, for a Fano manifold $X$ the nefness assumption on $T_{X}$ should imply the existence of an affine algebraic group acting transitively on $X$. One usually considers the case in which the Picard number $\rho_{X}$ is equal to one, from which the general case should follow. For more information on the Conjecture and further references we refer the reader to the survey [21].
In view of Mori’s proof of the Hartshorne–Fränkel Conjecture – of which the CP Conjecture is a natural generalization – one of the strategies that has been used to prove partial results is the study of rational curves. If the tangent bundle $T_{X}$ is nef, then the deformations of rational curves in $X$ are unobstructed, and one may try to reconstruct the geometry of $X$ out of certain families of rational curves. Of particular interest in this sense are the so-called unsplit and dominating families of rational curves, that allow us to consider some sort of infinitesimal-to-global arguments. More concretely, many results in the literature show that the homogeneity of a Fano manifold $X$ of Picard number one can be inferred from the homogeneity of the sets of tangent directions to the curves of an unsplit and dominating family of rational curves at points (see, for instance [20, 12, 27]).
@@
In order to prove Theorem 1.1 we first study, in Section 4, the induced ${\mathbb{C}}^{*}$-actions on the beautiful family ${\mathcal{M}}$ of rational curves in $X$ and on the universal family ${\mathcal{U}}$, without any assumptions on the extremal fixed point components. We prove that both the induced actions on ${\mathcal{M}}$ and ${\mathcal{U}}$ are equalized and we describe the induced action on the variety ${\mathcal{U}}_{y}$, defined as the inverse image of $y\in X$ by the evaluation morphism $q:{\mathcal{U}}\to X$.
@@
In Section 6 we first refine the above results in the case in which the action has an isolated extremal fixed point. In particular, if $y$ is a fixed point belonging to the fixed point component joined to the isolated point by curves in ${\mathcal{M}}$, the action on ${\mathcal{U}}_{y}$ has an extremal fixed point, too. Since the criticality of the action on ${\mathcal{U}}_{y}$ is low, these varieties can be classified: Section 3 is devoted to this task (see Theorems 3.6 and 3.9).
The proof of Theorem 1.1 then proceeds by discarding the non-homogeneous possibilities for ${\mathcal{U}}_{y}$ and by showing that, for every $x\in X$ the variety ${\mathcal{U}}_{x}$ is isomorphic to ${\mathcal{U}}_{y}$. The homogeneity of $X$ then follows from [27, Theorem 1.1], and the fact that $X$ is an irreducible Hermitian symmetric space from the classification of equalized ${\mathbb{C}}^{*}$-actions on homogeneous varieties (see [25, Table 2]).
@@ 2.1. ${\mathbb{C}}^{*}$-actions
$(T_{X})_{\mid Y}=T^{+}\oplus T^{0}\oplus T^{-},$ (1)
where $T^{+}$, $T^{0}$, $T^{-}$ are the subbundles of $(T_{X})_{\mid Y}$ where ${\mathbb{C}}^{*}$ acts with positive, zero or negative weights, respectively. Then, by local linearization, $T^{0}=T_{Y}$ and
@@
Throughout the paper we will freely use the notation $Y_{\pm}$, $N^{\pm}(Y)$, …, to refer to $Y_{+}$ and $Y_{-}$, $N^{+}(Y)$ and $N^{-}(Y)$, …
A straightforward consequence of the decomposition (2) is that the tangent bundle of a fixed point component is a quotient of the restriction of the tangent bundle of $X$.
@@ 2.2. Białynicki-Birula decomposition
We refer to [6] for a complete account on the Białynicki-Birula decomposition and its applications, and to [4] for the original reference. Given any set $S\subset X$, we define:
$X^{\pm}(S):=\{x\in X|\,\,x_{\pm}\in S\},\quad B^{\pm}(S):=\overline{X^{\pm}(S)}.$ (4)
@@
We will denote by $B^{\pm}(Y)$ the closure of the BB-cells $X^{\pm}(Y)$. Every cell $X^{\pm}(Y)$ has a natural morphism onto $Y$, sending every $x\in X^{\pm}(Y)$ to its limiting point $x_{\pm}$. It was shown by Białynicki-Birula that $X^{\pm}(Y)$ is an affine bundle, locally isomorphic to $N^{\pm}(Y)$.
@@ 2.3. Actions on polarized pairs
A ${\mathbb{C}}^{*}$-action on the polarized pair $(X,L)$ is a ${\mathbb{C}}^{*}$-action on $X$, together with the weight map $\mu_{L}$ determined by the normalized linearization of an ample line bundle $L$. In this case the minimum and maximum of the weight map $\mu_{L}$ are achieved precisely at the sink and the source of the action. We then denote by
@@
so that $Y_{0},Y_{r}$ are the sink $Y_{-}$ and the source $Y_{+}$ of the action on $X$, respectively.
@@ 2.4. Equalized actions
A ${\mathbb{C}}^{*}$-action is equalized at $Y\in{\mathcal{Y}}$ if for every $x\in\big{(}X^{-}(Y)\cup X^{+}(Y)\big{)}\setminus Y$ the isotropy group of the action at $x$ is trivial. Equivalently (see [23, Lemma 2.1]), ${\mathbb{C}}^{*}$ acts on $N^{+}(Y)$ with all the weights equal to $+1$ and on $N^{-}(Y)$ with all the weights equal to $-1$. The action is called equalized if it is equalized at every $Y\in{\mathcal{Y}}$.
For an equalized action the closure of any $1$-dimensional orbit is a smooth rational curve. In fact, if $x_{-}$ is the sink of a $1$-dimensional orbit ${\mathbb{C}}^{*}x$ and we denote by $Y\subset X$ the fixed point component containing $x_{-}$, then the BB-cell $X^{-}(Y)$ is locally ${\mathbb{C}}^{*}$-equivariantly isomorphic to $N^{-}(Y)$ around $x^{-}$. Since the action of ${\mathbb{C}}^{*}$ is equalized, the closures of the $1$-dimensional orbits of ${\mathbb{C}}^{*}$ in $N^{-}(Y)$ are lines in the fibers of $N^{-}(Y)\to Y$; in particular $\overline{{\mathbb{C}}^{*}x}$ is smooth at $x_{-}$; a similar argument applies to the source of the orbit.
The $L$-degree of the closure of a $1$-dimensional orbit may be computed in terms of the weights at its extremal points. (Cf. [23, Lemma 2.2], [28, Corollary 3.2]):
@@
Note that since every Cartier divisor on a projective variety $X$ is linearly equivalent to the difference of two ample divisors, we get the following:
@@
Let $X$ be a smooth projective variety with an equalized ${\mathbb{C}}^{*}$-action, $L$ a line bundle on $X$, and $C$ the closure of a $1$-dimensional orbit, with sink $y_{-}$ and source $y_{+}$. Then:
$L\cdot C=\mu_{L}(y_{+})-\mu_{L}(y_{-}).$
In particular, if the fixed point components containing the sink and the source of $C$ are denoted by $Y,Y^{\prime}$, respectively, we have that:
@@
Not even the equalization hypothesis is enough to guarantee that the BB-decompositions associated with a ${\mathbb{C}}^{*}$-action on a smooth projective variety are stratifications, see [24, Example 6.1].
@@ 2.5. Geometric quotients
As shown in [24, Section 2.2], $\mathcal{G}\!X(\tau)$ is the same for any $\tau\in(a_{i},a_{i+1})\cap{\mathbb{Q}}$, and so we set:
@@
If the ${\mathbb{C}}^{*}$-action on $X$ is equalized, the geometric quotients $\mathcal{G}\!X_{0,1}$ and $\mathcal{G}\!X_{r-1,r}$ can be described as follows (see [25, Section 2.1.3] for details): the ${\mathbb{C}}^{*}$-action on $X$ extends to an action on the blowup $X^{\flat}$ of $X$ along its sink and source. Then $\mathcal{G}\!X_{0,1}$ and $\mathcal{G}\!X_{r-1,r}$ are isomorphic to the sink and the source of $X^{\flat}$, that is to ${\mathbb{P}}(N_{Y_{0},X}^{\vee})$ and ${\mathbb{P}}(N_{Y_{r},X}^{\vee})$, respectively (see [25, Section 2.1.3]). In particular, in the case of an action of B-type, i.e., when the extremal fixed point components are divisorial (equivalently, if $\nu^{-}(Y_{-})=\nu^{+}(Y_{+})=1$), we have that $\mathcal{G}\!X_{0,1}\simeq\mathcal{G}\!X_{0,0}\simeq Y_{0}$ and $\mathcal{G}\!X_{r-1,r}\simeq\mathcal{G}\!X_{r,r}\simeq Y_{r}$.
The varieties $\mathcal{G}\!X_{i,i+1}$ are birationally equivalent, and the natural birational maps among them fit in the following commutative diagram:
@@
The diagonal maps between geometric quotients are contractions, the maps $s$ and $d$ are smooth blow-ups, the horizontal maps are Atiyah flips, and the squares are defined as fiber products (see [30, Theorem 1.9]).
In particular, in the case in which the action is of B-type, the diagram above provides a birational map $\psi:Y_{0}\dashrightarrow Y_{r}$.
The varieties $Z_{i,i+2}$ appearing in Diagram (7) have been interpreted in [24] in terms of Chow quotients of birational modifications of $X$ and the centers of the blowups $s$ and $d$ have been described. In this paper we will only use them as resolutions of the birational maps $\mathcal{G}\!X_{i,i+1}\dashrightarrow\mathcal{G}\!X_{i+1,i+2}$. More precisely, let us summarize in the next proposition what we will need later.
@@
The center of the blowup $s:Z_{i,i+2}\to\mathcal{G}\!X_{i,i+1}$ (resp. $d:Z_{i,i+2}\to\mathcal{G}\!X_{i+1,i+2}$) is the inverse image in $\mathcal{G}\!X_{i,i+1}$ (resp. $\mathcal{G}\!X_{i+1,i+2}$) of the embedding of $Y_{i+1}$ into $\mathcal{G}\!X_{i+1,i+1}$. Moreover, if $\nu^{-}(Y_{i+1})=1$ (resp. if $\nu^{+}(Y_{i+1})=1$) then $s$ (resp. $d$) is an isomorphism.
@@
The first part of the statement follows from [24, Remark 5.10], the second from [24, Proof of Corollary 7.1]. ∎
@@
Assume that we have a B-type action of criticality three on $X$ with divisorial sink and source $Y_{0},Y_{3}$ and that $\nu^{+}(Y_{1})=\nu^{-}(Y_{2})=1$. Then the birational map $\psi:Y_{0}\dashrightarrow Y_{3}$ is resolved by a smooth blowup with center isomorphic to $Y_{1}$ and a smooth blowdown with center isomorphic to $Y_{2}$.
@@ 2.6. Action on rational homogeneous varieties
Standard examples of ${\mathbb{C}}^{*}$-actions can be constructed easily on rational homogeneous (RH, for short) varieties, i.e., projective varieties that are quotients of a semisimple group $G$ by a parabolic subgroup $P$. Without loss of generality we assume that $G$ is the adjoint group of a semisimple Lie algebra ${\mathfrak{g}}$. Following [21, Section 2.2], an RH variety is completely determined by the Dynkin diagram ${\mathcal{D}}$ of $G$ marked on $\rho(G/P)$ nodes; we will then use the notation ${\mathcal{D}}(I)$ to denote the RH variety determined by the Dynkin diagram ${\mathcal{D}}$ marked in the set $I$ of nodes. Table 1 below illustrates how this notation works in the case of irreducible Hermitian symmetric spaces; these are the RH varieties $G/P$ of Picard number one such that $P$ acts irreducibly on the tangent space $T_{G/P,eP}$, which is $P$-equivariantly isomorphic to the quotient of the associated Lie algebras, ${\mathfrak{g}}/{\mathfrak{p}}$.
@@
The ${\mathbb{C}}^{*}$-action determined by $\sigma_{i}$ on the variety ${\mathcal{D}}(I)=G/P$ satisfies that $eP$ lies in the sink of the action. Furthermore, $\{eP\}$ is the isolated sink of the action if and only if $I=\{i\}$.
@@
Table 3. Irreducible Hermitian symmetric spaces and ${\mathbb{C}}^{*}$-actions with isolated sink on them.
@@ 2.7. Families of rational curves
Given a smooth complex projective variety $X$, a family of rational curves on $X$ is an irreducible component ${\mathcal{M}}$ of the normalization $\mathop{\rm RatCurves}\nolimits^{n}(X)$ of the scheme $\mathop{\rm RatCurves}\nolimits(X)$. Such a family comes equipped with a smooth ${\mathbb{P}}^{1}$-fibration $p:{\mathcal{U}}\to{\mathcal{M}}$, and an evaluation morphism $q:{\mathcal{U}}\to X$. The family ${\mathcal{M}}$ is called dominating if $q$ is dominant, and unsplit if ${\mathcal{M}}$ is proper. We will denote by ${\mathcal{U}}_{x}$ the fiber of $q$ over the point $x\in X$, and by ${\mathcal{M}}_{x}:=p(q^{-1}(x))\subset{\mathcal{M}}$ the family of rational curves of ${\mathcal{M}}$ passing by $x$.
In this paper we will consider the case in which ${\mathcal{M}}$ is a dominating, unsplit family such that $q:{\mathcal{U}}\to X$ is a smooth morphism, and we will simply say that the family is beautiful.
@@
By [17, II.Proposition 3.5], given an unsplit family of rational curves ${\mathcal{M}}$, it is beautiful if and only if the following property is fulfilled: $\nu^{*}T_{X}$ is globally generated – equivalently nef – for the normalization $\nu:{\mathbb{P}}^{1}\to X$ of any curve of the family. From this it follows that ${\mathcal{M}}$, ${\mathcal{U}}$ and ${\mathcal{M}}_{x}$, for every $x\in X$, are smooth, and that, for every element $C$ of ${\mathcal{M}}$:
$\dim{\mathcal{M}}_{x}=-K_{X}\cdot C-2.$ (8)
In particular, $\mathop{\rm RatCurves}\nolimits^{n}(X)$ is smooth at every point of ${\mathcal{M}}$, and every deformation of a curve of ${\mathcal{M}}$ belongs to ${\mathcal{M}}$. Then, given any irreducible subvariety $Y\subset X$, the set of elements of ${\mathcal{M}}$ parametrizing curves of the family contained in $Y$ is a union of irreducible components of $\mathop{\rm RatCurves}\nolimits^{n}(Y)$.
Remark 2.10.
Note that, if ${\mathcal{M}}$ is a beautiful family of rational curves in $X$ and $X$ is a Fano manifold (this will always be the case if the Picard number of $X$ is one), then every ${\mathcal{U}}_{x}$ is connected, therefore irreducible (cf. [15, Proposition 1.6]). Moreover, it is known that, via the morphism $p$, ${\mathcal{U}}_{x}\simeq{\mathcal{M}}_{x}$ for general $x\in X$ (cf. [21, Lemma 2.10]).
@@
Curves of ${\mathcal{M}}$ meeting $Z$ are parametrized by ${\mathcal{M}}_{Z}:=p(q^{-1}(Z))$, which is irreducible. If $q^{-1}(Z)\cap p^{-1}({\mathcal{M}}_{Z})=q^{-1}(Z)$ every curve of ${\mathcal{M}}$ meeting $Z$ would be contained in $Z$, but this is not possible since $X$ is ${\mathcal{M}}$)-connected. Therefore $q^{-1}(Z)\cap p^{-1}({\mathcal{M}}_{Z})$ is a proper closed subset of $p^{-1}({\mathcal{M}}_{Z})$, and so it does not contain a general fiber of $p^{-1}({\mathcal{M}}_{Z})\to{\mathcal{M}}_{Z}$. ∎
@@
Let $X$ be a Fano manifold of Picard number one, and let $p:{\mathcal{U}}\to{\mathcal{M}}$ be a beautiful family of rational curves in $X$. If the fiber ${\mathcal{U}}_{x}=q^{-1}(x)$ is a rational homogeneous variety for every $x\in X$, then $X$ is rational homogeneous.
@@
In the spirit of the above statement, irreducible Hermitian symmetric spaces can be characterized by the homogeneous varieties ${\mathcal{U}}_{x}=q^{-1}(x)$, which in this case is isomorphic to the variety ${\mathcal{M}}_{x}$ parametrizing lines passing by $x$, for every $x$. Table 4 describes ${\mathcal{M}}_{x}$ for every type of irreducible Hermitian symmetric space.
${\mathcal{D}}(I)$ ${\mathcal{M}}_{x}\simeq{\mathcal{U}}_{x}$
@@
Note that, by comparing this with Table 3, we immediately see that, in each case, ${\mathcal{M}}_{x}$ coincides with the component $Y_{1}$ of fixed points of weight one of the corresponding ${\mathbb{C}}^{*}$-action with isolated sink.